OpenStudy (anonymous):

Multiplying a Trinomial and Binomial: (k² - 4k + 3) (k - 2) Please explain

OpenStudy (anonymous):

What you need to do is use the Distributive Property to solve. Take each term in the first set of parenthesis, and distribute it to each term in the second set of parenthesis. For this question, I will give you the answer so you can use it as an example for next time to have a question like this: \[k ^{3} - 6k ^{2} + 11k - 6\]

OpenStudy (anonymous):

*you

OpenStudy (jdoe0001):

the same as if you were multiplying a binomial and a binomial :) you distribute the factors \(\large (k^2 - 4k + 3) (k - 2)\implies \begin{array}{llll} k^2 - 4k + 3\\ \qquad \times k\\ \hline \end{array}+\begin{array}{llll} k^2 - 4k + 3\\ \qquad \times -2\\ \hline \end{array}\)

OpenStudy (anonymous):

I get it, thank you! @iamgrace and @jdoe0001

OpenStudy (jdoe0001):

yw